Solve the equation:
x/4 -1= x/5
I dont understand at all how to do it.

Answer by dolly(163)

x/4 -1= x/5
The LCD of the denominators = 20
Multiplying by the LCD, throughout,
20(x/4) - 20(1) = 20(x/5)
==> 5x - 20 = 4x
==> 5x - 4x = 20
==> x = 20